# Exploring the Privacy-Energy Consumption Tradeoff for Split Federated Learning

### 05 Jan 2024 v2

@@ Abstract
Split Federated Learning (SFL) has recently emerged as a promising distributed learning technology, leveraging the strengths of both federated learning and split learning. It emphasizes the advantages of rapid convergence while addressing privacy concerns. As a result, this innovation has received significant attention from both industry and academia. However, since the model is split at a specific layer, known as a cut layer, into both client-side and server-side models for the SFL, the choice of the cut layer in SFL can have a substantial impact on the energy consumption of clients and their privacy, as it influences the training burden and the output of the client-side models. Moreover, the design challenge of determining the cut layer is highly intricate, primarily due to the inherent heterogeneity in the computing and networking capabilities of clients. In this article, we provide a comprehensive overview of the SFL process and conduct a thorough analysis of energy consumption and privacy. This analysis takes into account the influence of various system parameters on the cut layer selection strategy. Additionally, we provide an illustrative example of the cut layer selection, aiming to minimize the risk of clients from reconstructing the raw data at the server while sustaining energy consumption within the required energy budget, which involve trade-offs. Finally, we address open challenges in this field. These directions represent promising avenues for future research and development.
1 Introduction
@@
Federated Learning (FL) fundamentally addresses the challenges associated with centralized learning by distributing the training process across multiple clients, enabling parallel processing. This approach also helps safeguard the privacy of raw data stored on clients by exchanging only model parameters. However, FL necessitates local training on each client, which can be a significant burden on clients with limited battery power and computational resources when dealing with large models like Deep Learning (DL). To mitigate this problem, Split Learning (SL) has emerged as a solution. SL involves breaking down a full DL model into two sub-models, which can be trained both at a main server and across distributed clients. This approach alleviates the local training burden associated with FL while preserving data privacy. Nevertheless, SL introduces its own set of challenges, primarily related to training time overhead, owing to its relay-based training method. In this relay-based approach, only one client engages in the training process with the main server at any given time, while other clients remain in an idle state. This sequential training method leads to inefficient distributed processing, resulting in long training latency. To address this challenge, various strategies for parallelizing the training process of SL have been introduced [1]. Inspired by these efforts, Split Federated Learning, simply called SplitFed Learning, (SFL) has been recently proposed as a novel approach that leverages the strengths of both FL and SL. Unlike SL, in SFL, all clients perform their local training in parallel while actively engaging with the main server and the federated server (fed server). In SFL, the fed server plays a pivotal role in aggregating the local model updates from clients using pre-defined aggregation techniques, such as FedAvg [^1] FedAvg is the most popularly used weighted aggregation in FL, assigning varying weights to clients during aggregation based on the size of their datasets. This aggregation process occurs synchronously in each round of training. By introducing this additional aggregation server, SFL combines the advantages of both FL and SL seamlessly [2].
Despite the advantageous integration of SL and FL in SFL, the current SFL still presents several privacy concerns. Notably, the exchanged output of the client-side models, known as smashed data, and model updates between clients and servers have correlations with the raw data, leaving them susceptible to potential reconstruction attacks, where adversaries attempt to reconstruct the raw data from the correlated information. In this regard, differential privacy (DP) is currently the gold standard for privacy, known for its effectiveness against reconstruction and membership inference attacks by adding noise to the data or query responses that are released to untrusted parties, thereby obfuscating the outputs related to the sensitive data [3, 4]. Nevertheless, the adoption of DP mechanisms presents an added computational load for clients. Furthermore, since SFL still imposes notable latency due to computation and communication overhead until the model converges, implementing sophisticated DP can also pose practical deployment challenges.
Figure 1: Workflow of SFL: Combining both FL and SL perspectives.
@@
To tackle this growing concern, this article advocates for determining an appropriate split point between client-side and server-side models, specifically known as cut layer selection in the SFL, with the goal of mitigating the risk of reconstruction attacks while sustaining energy consumption within the required energy budget. This is particularly important for clients operating on limited battery power. It is worth noting that the selection of the cut layer significantly influences the outcome of the smashed data, impacting computational and communication overhead, as well as the level of privacy. For example, as discussed in [5], there is an ongoing study focused on the adaptive selection of cut layer, taking into account the varying computing and networking capabilities of clients. Nevertheless, prior research has not thoroughly analyzed both energy consumption and privacy in the context of cut layer selection, despite some studies focusing on empirical and analytical studies of reducing overall latency in SFL. This motivates us to provide an analysis of both energy consumption and privacy levels related to cut layer selection and a potential solution for developing novel cut layer selection techniques.
@@
2 Why is Cut Layer Selection Important?
2.1 Background about SFL
@@
The SFL framework consists of two servers such as i) a fed server and ii) main server, and multiple clients. The entire model is divided into two distinct sub-models: i) the client-side model and ii) the server-side model. As a practical concern, to reduce communication latency, the main server and fed server are set up using either of containers (such as Kubernetes) or virtual machines on the same or different physical multi-access computing (MEC) servers. The workflow of SFL is depicted in Fig. 1. Each client performs forward propagation on the client-side model using its dataset and passes the smashed data and corresponding label to the centralized main server (Step 1-2). The main server performs forward propagation from the smashed data on the server-side model and back-propagation by calculating the loss between the true label and predicted label, which can be done in parallel. Afterward, the server-side model is obtained by using FedAvg (Step 3-4) and the main server passes the gradient of the smashed data to each client for the client-side local model update (Step 5-6). Then, the fed server receives updated client-side local models from all clients and aggregates them by using FedAvg (Step 7-8). Finally, the fed server sends the updated client-side global model to all clients, enabling synchronization of client-side models (Step 9).
Figure 2: Example of reconstruction attacks in SFL as a privacy concern: This highlights the tradeoff between the client model complexity and privacy. Additionally, it explores the effects of Gaussian noise addition on privacy issues.
2.2 A Brief Survey of SFL
2.2.1 Communication efficiency Issues
@@
Despite the advantageous integration of SL and FL in the SFL to alleviate the computational burden on clients and address privacy concerns, it still imposes significant communication overhead. This is primarily due to the transmission of smashed data and gradients between clients, the main server, and the fed server, indicating the need for improvements [6]. Nevertheless, there remains a scarcity of research focusing on the analytical modeling of SFL, in contrast to the abundance of empirical studies [7]. Notably, [8, 2] have contributed analytical models that investigate the impact of total model training time (i.e., latency) concerning the cut layer point. Correspondingly, based on such analysis, there have been various approaches to improve the communication efficiency in SFL [6, 8, 1]. Specifically, [8] introduced an optimal cut layer selection based on latency analysis. Remarkably, most prior research has concentrated on analyzing and optimizing latency in SFL, yet there is a notable absence of studies addressing the impact of energy consumption concerning the cut layer selection in SFL. This aspect is crucial for estimating the associated costs for clients with limited battery power.
2.2.2 Privacy Issues
@@
In SFL, privacy concerns arise from the interactions between clients and servers, including both the main server and the fed server. During the forward phase, when clients transmit smashed data to the main server, this data becomes susceptible to reconstruction attacks [4], jeopardizing the privacy of the original information. Furthermore, the exchange of model updates with the fed server introduces another potential vulnerability to the raw data. To mitigate these risks, several privacy-preserving mechanisms can be employed. One effective strategy is the implementation of noise addition, where random noise is added to the data or model updates, thereby disrupting any attempts to reconstruct the original data. This method, often used in conjunction with subsampling techniques, ensures that each update’s information is based on a random subset of the data, further complicating potential attacks. Another approach is homomorphic encryption, which allows computations to be performed on encrypted data, enabling the server to process data updates without actually viewing the original data. These mechanisms, rooted in the principles of DP, work synergistically to protect client data during both the forward and backward phases of interaction, thereby reinforcing the confidentiality of sensitive information amidst the continual exchanges in federated environments. However, it is worth noting that this approach does introduce added complexity on the client side. In a recent study [9] on SL, researchers conducted an empirical investigation into the impact of selecting the cut layer on reconstruction attacks during the forward phase. Their findings indicate that a greater depth of layers on the client side implies more non-linear functions, which compress the raw data by eliminating less informative features, thereby enhancing its resistance to reconstruction attacks. Similarly, the work of [10] also presented empirical studies on how the distance correlation between raw data and smashed data influences privacy leakage in SL in relation to the selection of cut layer. Building on this insight, we will conduct empirical studies of SFL in a similar fashion, which can be extended to explore cut layer selection strategies that strike a balance between energy consumption and privacy level.
@@
2.3 Cut Layer Selection Impact
2.3.1 Energy Efficiency Perspective
@@
2.3.2 Privacy Level Perspective
@@
As shown in Fig. 2 , when we assume that the main server is honest but curious (HBC), it is still possible for the main server to reconstruct the original input data from the smashed data (specifically, the embedded features) received from clients. This process is referred to as an reconstruction attack. In this context, we can assess the privacy risks associated with this reconstruction attack by measuring the human perceptual similarity between the original and the reconstructed images, known as the structural similarity index (SSIM). This SSIM, which ranges from [0, 1], is widely used as a metric to measure the attack and defense performance, where 1 denotes the most similar. Fig. 2 illustrates the relationship between client model complexity and privacy leakage. It highlights the effects of altering the cut layer on the quality of images reconstructed at the main server. Notably, as the cut layer index increases, indicating a rise in model complexity, the quality of the reconstructed image degrades. This degradation is interchangeable with enhanced privacy. The reason is that the complexity of the model deployed at the client’s end introduces advanced non-linearities in its output. This added intricacy can make it more challenging for adversaries to reverse-engineer and retrieve private input data. Moreover, Fig. 2 also presents an illustrative example to practically demonstrate the application of noise addition, a fundamental strategy in achieving differential privacy. Here, our focus will center on one widely adopted noise addition mechanism: the Gaussian mechanism where it is instrumental in introducing carefully calibrated noise to sensitive data, effectively masking individual information while preserving the utility of the overall dataset. The Gaussian mechanism employs a Gaussian distribution for this purpose, providing distinct privacy-accuracy trade-offs. In this context, it is evident that introducing Gaussian noise to smashed data further degrades the quality of the reconstructed image, as indicated by the mean square error (MSE) representing the differences in pixels between the original and the reconstructed images. As the level of Gaussian noise rises, the MSE also increases, leading to a decrease in the SSIM in both shallow and deep cut layer scenarios. We also confirmed that the classification accuracy remains within an acceptable range (i.e., 92.6-94.4%) regardless of the level of Gaussian noise, which means that Gaussian mechanism has less sacrifice of accuracy while having larger privacy gains. This simple example will serve to illuminate the intricacies and operational dynamics of this noise addition technique, offering insights into their strategic implementation in contexts demanding stringent data privacy measures [^2] Various noise addition mechanisms exist, including the Laplacian mechanism, which introduces noise following a Laplace distribution. However, since the selection of noise addition mechanisms is beyond the scope of this paper, we focus on exploring the impact of Gaussian noise to provide straightforward insight..
3 Case Study: Cut Layer Selection
3.1 An Illustrative Example
@@
Based on the latency analysis of SFL in [2] and [8], and leveraging the energy consumption insights gained from FL studies [11], this section examines the overall energy consumption associated with SFL. Note that the analysis is based on the vanilla version of SFL rather than SFLv2/v3. However, the core concepts and intuitions are easily applicable to the higher versions. Similar to FL, as shown in Fig. 3, the total energy consumption $E$ of each client during SFL encompasses both i) computing and ii) networking energy. However, it requires customization to account for model splitting and interactions with both the fed server and the main server, distinctive features introduced in the SFL. Taking these considerations into account, to formulate the energy consumption model, consider a scenario with $K$ clients. Then, $K$ clients engage in the SFL framework for $g_{e}$ global epochs, representing the overall number of rounds necessary to achieve a specific training loss [12]. The client-side models synchronize at each global epoch, enabled by the fed server. In each global epoch, leveraging this synchronized client-side model, both clients and the main server conduct training for $l_{e}$ local epochs. Specifically, for simplicity, as in [2] and [8], considering homogeneous clients in a single local epoch, each client trains $D_{b}$ randomly sampled data items, commonly referred to as a mini-batch, from its dataset. By using transmission power $P_{t}$, they can communicate with the fed server and main server with total uplink (downlink) transmission rates $R_{1}^{U}$ and $R_{2}^{U}$ ($R_{1}^{D}$ and $R_{2}^{D}$), respectively. Consequently, the uplink (downlink) transmission rates for each client reduce to $\frac{R_{1}^{U}}{K}$ and $\frac{R_{2}^{U}}{K}$ ($\frac{R_{1}^{D}}{K}$ and $\frac{R_{2}^{D}}{K}$), respectively. Note that the entire model $W=[W_{C};W_{S}],$ where $W_{C}$, and $W_{S}$ are the client-side model, and the server-side model, respectively. As in [2] and [8], we also assume the same size of smashed data denoted as $q$ sent from clients to the main server. Let $|W|$ be the number of model parameters in the entire model $W$ where all model parameters have the same size $b$, and let $\alpha$ be the fraction of model parameters, which serves as the model cut layer point for SFL in $W_{C}$, where $|W_{C}|=\alpha|W|$ and $|W_{S}|=(1-\alpha)|W|$. And, $T$ represents the time taken for one forward and backward propagation on the full model. In backward propagation, gradients of a constant size, denoted as $q^{\prime}$, are transmitted. Therefore, in SFL, energy consumption during interaction with the main server, denoted as $E_{main}$, requires each client i) to proceed forward and backward propagation using $W_{C}$, incurring energy consumption $\alpha TP_{c}$ and ii) to transmit the smashed data (or receive the gradients) to (from) the main server per each sampled data item, incurring energy consumption $\frac{qK}{R_{2}^{U}}P_{t}$ (or $\frac{q^{\prime}K}{R_{2}^{D}}P_{r}$). Here, $P_{r}$ is the receiving power consumption at each client. In parallel, energy consumption for interaction with the fed server, denoted as $E_{fed}$, requires each client i) to send its local model to the fed server aggregation, incurring energy consumption $\frac{\alpha b|W|K}{R_{1}^{U}}P_{t}$, and ii) to receive the updated global model from the fed server, incurring energy consumption $\frac{\alpha b|W|K}{R_{1}^{D}}P_{r}$ during one global epoch. Finally, total energy consumption $E(\alpha)$ of each client for the SFL is given by
$E(\alpha)=g_{e}(l_{e}D_{b}E_{main}+E_{fed}).$ (1)
@@
As previously discussed, under the assumption of an HBC main server, the privacy risks associated with the reconstruction attack can be evaluated by measuring the SSIM between the original input and the reconstructed images with respect to $\alpha$, denoted as $RS(\alpha)$. To evaluate the impact of cut layer selection in terms of $RS(\alpha)$ and $E(\alpha)$, we utilize a training-based adversarial inversion approach [13]. Note that the reconstruction model is implemented to reflect the inverted structure of classifier $W$ with transposed convolutional and Tanh activation layers. In our experiment, the overall classifier model $W$ for SFL is designed with four convolution blocks (Conv-BN-ReLU-Conv-BN-ReLU-Conv-BN-Max-ReLU) and one linear block. Linear block consists of two fully-connected layers and a Softmax function. The cut layer is selected between blocks to divide client and server-side models. For performance evaluation, the Fashion-MNIST dataset was employed [^3] https://github.com/zalandoresearch/fashion-mnist. The input image was resized to $32\times 32$ and the $D_{b}$ was 128. We used an Adam optimizer and the learning rate was set to 0.0002. The betas for Adam were set to (0.5, 0.999). We trained the classifier model for SFL up to 50 global training epochs and 75 local training epochs. The reconstruction model was trained up to 50 epochs under the same optimizer and settings as the classifier model. More detailed parameters are summarized at Table I.
@@
As depicted in Fig. 4, the RS exhibits a consistent decrease as the depth of the cut layer increases. This phenomenon implies that the depth of the cut layer can be directly correlated with the complexity of the client-side model, potentially make attackers to reconstruct the image from the smashed data more difficult. Consequently, in this empirical study, we demonstrated that a convex model can effectively approximate the behavior of RS, yielding a Root Mean Square Error (RMSE) of only $0.0019$. As a result, RS can be estimated as $RS(\alpha)$ = $0.76860.367\alpha^{2}-0.7045\alpha+$ 0.7686. Consequently, we can optimize the selection of the cut layer, denoted by $\alpha$, to minimize privacy leakage from reconstruction attacks, as measured by $RS(\alpha)$, while ensuring that energy consumption ($E$) remains within an acceptable energy budget ($E_{req}$). Thus, the problem can be formulated as follows:
@@
Note that $E(\alpha)$ within the constraint should be calculated as an average across clients when dealing with heterogeneous client scenarios. The problem of (2) is a convex optimization problem since it features a convex objective function and affine constraints, which guarantees global optimal solution and can be simply solved using a CVX solver. Based on our empirical findings, the objective function $RS(\alpha)$ is strictly decreasing. Therefore, the optimal cut layer selection is expected to be located at the constraint boundary. In practice, a continuous value for $\alpha$ should be translated into an integer-valued cut layer index. If the depth of the entire network is not excessively large and the estimated $RS(\alpha)$ and $E(\alpha)$ function are not provided, exhaustive search remains a viable option. Based on the intuition from analytical study, this exhaustive search involves verifying the energy budget constraint as the depth of the cut layer increases. Thus, as depicted in Fig. 5, by optimizing the cut layer selection while ensuring the energy consumption budget, $R(\alpha)$ can be minimized in comparison to other cut layer selections within the feasible range of $\alpha$. This indicates that a thoughtful cut layer selection can strike a balance between privacy levels and energy consumption. In practical scenarios, such cut layer selection can be jointly optimized with other system parameters, all working towards similar objectives as well as multi-objectives (i.e., including latency). This may result in a more intricate problem, which we will delve into in the following subsection.
@@
Figure 5: Privacy gain of proposed scheme: It represents that optimal cut layer selection achieves minimum privacy leakage while satisfying the required energy budget.
3.2 Other Types of Control Variables

### 20 Mar 2024 v3

@@ Abstract
Split Federated Learning (SFL) has recently emerged as a promising distributed learning technology, leveraging the strengths of both federated and split learning. It emphasizes the advantages of rapid convergence while addressing privacy concerns. As a result, this innovation has received significant attention from both industry and academia. However, since the model is split at a specific layer, known as a cut layer, into both client-side and server-side models for the SFL, the choice of the cut layer in SFL can have a substantial impact on the energy consumption of clients and their privacy, as it influences the training burden and the output of the client-side models. In this article, we provide a comprehensive overview of the SFL process and thoroughly analyze energy consumption and privacy. This analysis considers the influence of various system parameters on the cut layer selection strategy. Additionally, we provide an illustrative example of the cut layer selection, aiming to minimize clients’ risk of reconstructing the raw data at the server while sustaining energy consumption within the required energy budget, which involves trade-offs. Finally, we address open challenges in this field. These directions represent promising avenues for future research and development.
I Introduction
@@
Federated Learning (FL) fundamentally addresses the challenges associated with centralized learning by distributing the training process across multiple clients, enabling parallel processing. This approach also helps safeguard the privacy of raw data stored on clients by exchanging only model parameters. However, FL necessitates local training for each client, which can significantly burden clients with limited battery power and computational resources when dealing with large models like Deep Learning (DL). Split Learning (SL) has emerged as a solution to mitigate this problem. SL involves breaking down a full DL model into two sub-models, which can be trained both at a main server and across distributed clients. This approach alleviates the local training burden associated with FL while preserving data privacy. Nevertheless, SL introduces its own set of challenges, primarily related to training time overhead, owing to its relay-based training method. In this relay-based approach, only one client trains with the main server at any given time while other clients remain idle. This sequential training method leads to inefficient distributed processing and long training latency. To address this challenge, various strategies for parallelizing the training process of SL have been introduced [1]. Inspired by these efforts, Split Federated Learning, simply called SplitFed Learning (SFL), has been recently proposed as a novel approach that leverages the strengths of both FL and SL. Unlike SL, in SFL, all clients perform their local training in parallel while actively engaging with the main server and the federated server (fed server). In SFL, the fed server plays a pivotal role in aggregating the local model updates from clients using pre-defined aggregation techniques, such as FedAvg. This aggregation process occurs synchronously in each round of training. By introducing this additional aggregation server, SFL combines the advantages of both FL and SL seamlessly [2].
Figure 1: Workflow of SFL: In FL, the full model undergoes local training rather than just a subset. In SL, training processes do not occur in parallel. SFL seamlessly integrates the benefits of both FL and SL by addressing their respective limitations.
@@
Despite the advantageous integration of SL and FL in SFL, the current SFL still presents several privacy concerns. Notably, the exchanged output of the client-side models, known as smashed data, and model updates between clients and servers have correlations with the raw data, leaving them susceptible to potential reconstruction attacks, where adversaries attempt to reconstruct the raw data from the correlated information. To tackle this growing concern, this article advocates for determining an appropriate split point between client-side and server-side models, specifically known as cut layer selection in the SFL, to mitigate the risk of reconstruction attacks while sustaining energy consumption within the required energy budget. This is particularly important for clients operating on limited battery power. It is worth noting that the selection of the cut layer significantly influences the outcome of the smashed data, impacting computational and communication overhead, as well as the level of privacy. For example, as discussed in [3], an ongoing study focused on the adaptive selection of cut layer, taking into account the varying computing and networking capabilities of clients. Nevertheless, prior research has not thoroughly analyzed both energy consumption and privacy in the context of cut layer selection, despite some studies focusing on empirical and analytical studies of reducing overall latency in SFL. This motivates us to provide an analysis of both energy consumption and privacy levels related to cut layer selection and a potential solution for developing novel cut layer selection techniques.
@@
II Why is Cut Layer Selection Important?
II-A Background about SFL
@@
The SFL framework consists of two servers such as i) a fed server and ii) a main server and multiple clients. The entire model is divided into two distinct sub-models: i) the client-side model and ii) the server-side model. The workflow of SFL is depicted in Fig. 1. Each client performs forward propagation on the client-side model using its dataset, passing the smashed data and corresponding label to the centralized main server (Step 1-2). Notably, the smashed data is the features extracted from its dataset via forward propagation using the client-side model. The main server performs forward propagation from the smashed data on the server-side model and back-propagation by calculating the loss between the true label and the predicted label, which can be done in parallel. Afterward, the server-side model is obtained by using FedAvg (Step 3-4) and the main server passes the gradient of the smashed data to each client for the client-side local model update (Step 5-6). Then, the fed server receives all clients’ updated client-side local models and aggregates them using FedAvg (Step 7-8). Finally, the fed server sends all clients the updated client-side global model, enabling synchronization of client-side models (Step 9). In summary, in the SFL approach, clients are responsible for training only a part of the model, which reduces their computational burden compared to traditional FL. Furthermore, SFL enhances the convergence speeds of SL by synchronizing client-side models through the fed server. Nonetheless, interacting with two servers in SFL introduces an additional communication overhead.
Figure 2: Example of reconstruction and membership inference attacks in SFL as a privacy concern: This highlights the tradeoff between the client model complexity and privacy. Note that A reconstruction attack is an attempt to reconstruct the original input data from the smashed data. Additionally, membership inference attacks are another type of attack, aiming to determine whether a given data record was part of the target model’s training dataset or not. (The random forest-based attack model is commonly employed in such scenarios.)
II-B A Brief Survey of SFL
@@
As we discussed, SFL imposes significant communication overhead. This is primarily due to the transmission of smashed data, gradients and model updates between clients, the main server, and the fed server, indicating the need for improvements [4]. Nevertheless, there remains a scarcity of research focusing on the analytical modeling of SFL, in contrast to the abundance of empirical studies [5]. Notably, [6, 2] have contributed analytical models that investigate the impact of total model training time (i.e., latency) concerning the cut layer point. Correspondingly, based on such analysis, there have been various approaches to improve communication efficiency in SFL [4, 6, 1]. Specifically, [6] introduced an optimal cut layer selection based on latency analysis.
@@
Moreover, In SFL, privacy concerns arise from the interactions between clients and servers, including the main and fed servers. During the forward phase, when clients transmit smashed data to the main server, this data becomes susceptible to reconstruction attacks , jeopardizing the privacy of the original information. Furthermore, the exchange of model updates with the fed server introduces another potential vulnerability to the raw data. To mitigate these risks, several privacy-preserving mechanisms can be employed. In a recent study [7] on SL, researchers conducted an empirical investigation into the impact of selecting the cut layer on reconstruction attacks during the forward phase. Their findings indicate that a greater depth of layers on the client side implies more non-linear functions, which compress the raw data by eliminating less informative features, thereby enhancing its resistance to reconstruction attacks. Similarly, the work of [8] also presented empirical studies on how the distance correlation between raw data and smashed data influences privacy leakage in SL about the selection of cut layer. Remarkably, most prior research has concentrated on analyzing and optimizing latency in SFL. Yet, there is a notable absence of studies addressing the impact of energy consumption concerning the cut layer selection in SFL. This aspect is crucial for estimating the associated costs for clients with limited battery power. Moreover, building on the insight of SL regarding privacy concerns, we will similarly conduct empirical studies of SFL, which can be extended to explore cut layer selection strategies that strike a balance between energy consumption and privacy level.
@@
II-C Cut Layer Selection Impact
II-C1 Energy Efficiency Perspective
@@
II-C2 Privacy Level Perspective
@@
As shown in Fig. 2 , when we assume that the main server is honest but curious (HBC), it is still possible for the main server to reconstruct the original input data from the smashed data (specifically, the embedded features) received from clients. This process is referred to as a reconstruction attack. In this context, we can assess the privacy risks associated with this reconstruction attack by measuring the human perceptual similarity between the original and the reconstructed images, known as the structural similarity index (SSIM) [^1] This study concentrates on image data in the reconstruction attack, a prevalent focus in this area. Recently, the significance of text data privacy in this attack has also been recognized. In this case, metrics such as recovery rate, fluency, and attack accuracy are used to evaluate the privacy risk.. The SSIM formula compares two images by measuring three factors: luminance, contrast, and structure. It calculates the comparison as $\operatorname{SSIM}(x,y)=l(x,y)c(x,y)s(x,y)$, where $x$ and $y$ are the pixel values of the images being compared. The luminance comparison is denoted by $l$, contrast comparison by $c$, and structural comparison by $s$. This metric, which ranges from [0, 1], is widely used as a metric to measure the attack and defense performance [9], where the value $1$ denotes the most similar and vice versa. Fig. 2 (a)-(d) illustrates the relationship between client model complexity and privacy leakage. It highlights the effects of altering the cut layer on the quality of images reconstructed at the main server. Notably, as the cut layer index increases, indicating a rise in model complexity, the quality of the reconstructed image degrades. This degradation is interchangeable with enhanced privacy. The reason is that the complexity of the model deployed at the client’s end introduces advanced non-linearities in its output. This added intricacy can make it more challenging for adversaries to reverse-engineer and retrieve private input data. Moreover, Fig. 2 (a)-(d) also presents an illustrative example to practically demonstrate the application of noise addition, a fundamental strategy in achieving differential privacy. Here, our focus will center on one widely adopted noise addition mechanism: the Gaussian mechanism where it is instrumental in introducing carefully calibrated noise to sensitive data, effectively masking individual information while preserving the utility of the overall dataset. The Gaussian mechanism employs a Gaussian distribution for this purpose, providing distinct privacy-accuracy trade-offs. In this context, it is evident that introducing Gaussian noise to smashed data further degrades the quality of the reconstructed image, as indicated by the mean square error (MSE) representing the differences in pixels between the original and the reconstructed images. As the level of Gaussian noise rises, the MSE also increases, leading to a decrease in the SSIM in both shallow and deep cut layer scenarios. We also conduct another privacy assessment of the Membership Inference Attack (MIA), which is another type of attack on data privacy: determine whether a given data record was part of the target model’s training dataset or not. Here, the main server can still exploit these vulnerabilities to train a binary classifier and infer the membership information by taking the predicted value of the data sample obtained by querying the target model as input to the binary attack model. In this context, we can assess the privacy risks associated with this MIA by measuring the accuracy, which was widely used as a metric in MIA; the percentage of data records correctly inferred by the attack model in the whole given dataset. To train the attack model, we use transfer-inherit shadow learning. Fig. 2 (e) illustrates that MIA accuracy decreases with deeper cut layers and higher noise levels. Cutting at a shallow layer results in the main server retaining significant information about the entire target model, requiring less effort in the creation of a shadow client model. This makes the SFL vulnerable to MIA attacks, while the attack accuracy becomes more resilient to noise.
III Case Study: Cut Layer Selection
III-A An Illustrative Example
@@
Based on the latency analysis of SFL in [2] and [6], and leveraging the energy consumption insights gained from FL studies [10], this section examines the overall energy consumption associated with SFL. Note that the analysis is based on the vanilla version of SFL rather than SFLv2/v3. However, the core concepts and intuitions are easily applicable to the higher versions. Similar to FL, as shown in Fig. 3, the total energy consumption $E$ of each client during SFL encompasses both i) computing and ii) networking energy. However, it requires customization to account for model splitting and interactions with both the fed server and the main server, distinctive features introduced in the SFL. Taking these considerations into account, to formulate the energy consumption model, consider a scenario with $K$ clients. Then, $K$ clients engage in the SFL framework for $g_{e}$ global epochs, representing the overall number of rounds necessary to achieve a specific training loss [11]. The client-side models synchronize at each global epoch, enabled by the fed server. In each global epoch, leveraging this synchronized client-side model, both clients and the main server conduct training for $l_{e}$ local epochs. Specifically, for simplicity, as in [2] and [6], considering homogeneous clients in a single local epoch, each client trains $D_{b}$ randomly sampled data items, commonly referred to as a mini-batch, from its dataset. By using transmission power $P_{t}$, they can communicate with the fed server and main server with total uplink (downlink) transmission rates $R_{1}^{U}$ and $R_{2}^{U}$ ($R_{1}^{D}$ and $R_{2}^{D}$), respectively. Consequently, the uplink (downlink) transmission rates for each client reduce to $\frac{R_{1}^{U}}{K}$ and $\frac{R_{2}^{U}}{K}$ ($\frac{R_{1}^{D}}{K}$ and $\frac{R_{2}^{D}}{K}$), respectively. Note that the entire model $W=[W_{C};W_{S}],$ where $W_{C}$, and $W_{S}$ are the client-side model, and the server-side model, respectively. As in [2] and [6], we also assume the same size of smashed data denoted as $q$ sent from clients to the main server. Let $|W|$ be the number of model parameters in the entire model $W$ where all model parameters have the same size $b$, and let $\alpha$ be the fraction of model parameters, which serves as the model cut layer point for SFL in $W_{C}$, where $|W_{C}|=\alpha|W|$ and $|W_{S}|=(1-\alpha)|W|$ [^2] As the depth of a DL model increases, there is a corresponding rise in both the number of learnable parameters and multiply-add floating-point operations [12]. Conversely, SFL divides the layers of a model and assigns only a fraction determined by $\alpha$ to a client, indicating that lower $\alpha$ values correspond to reduced client-side model complexity.. And, $T$ represents the time taken for one forward and backward propagation on the full model. In backward propagation, gradients of a constant size, denoted as $q^{\prime}$, are transmitted. Here, $P_{c}$ represents the power consumption necessary to train the full model $W$ at the client, which is calculated as the total energy consumption required to train the entire model, divided by the duration $T$. Therefore, in SFL, energy consumption during interaction with the main server, denoted as $E_{\text{main}}$, requires each client i) to proceed forward and backward propagation using $W_{C}$, incurring energy consumption $\alpha TP_{c}$ and ii) to transmit the smashed data (or receive the gradients) to (from) the main server per each sampled data item, incurring energy consumption $\frac{qK}{R_{2}^{U}}P_{t}$ (or $\frac{q^{\prime}K}{R_{2}^{D}}P_{r}$). Here, $P_{r}$ is the receiving power consumption at each client. In parallel, energy consumption for interaction with the fed server, denoted as $E_{\text{fed}}$, requires each client i) to send its local model to the fed server aggregation, incurring energy consumption $\frac{\alpha b|W|K}{R_{1}^{U}}P_{t}$, and ii) to receive the updated global model from the fed server, incurring energy consumption $\frac{\alpha b|W|K}{R_{1}^{D}}P_{r}$ during one global epoch. Finally, total energy consumption $E(\alpha)$ of each client for the SFL is given by
$E(\alpha)=g_{e}(l_{e}D_{b}E_{\text{main}}+E_{\text{fed}}).$ (1)
@@
As previously discussed, under the assumption of an HBC main server, the privacy risks associated with the reconstruction attack can be evaluated by measuring the SSIM between the original input and the reconstructed images with respect to $\alpha$, denoted as $RS(\alpha)$. To evaluate the impact of cut layer selection in terms of $RS(\alpha)$ and $E(\alpha)$, we utilize a training-based adversarial inversion approach [13]. Note that the reconstruction model is implemented to reflect the inverted structure of the client-side model $W_{C}$ with transposed convolutional and Tanh activation layers. In our experiment, the overall classifier model $W$ for SFL is designed with four convolution blocks (Conv-BN-ReLU-Conv-BN-ReLU-Conv-BN-Max-ReLU) and one linear block. Linear block consists of two fully-connected layers and a Softmax function. The cut layer is selected between blocks to divide client and server-side models. For performance evaluation, the Fashion-MNIST dataset was employed [^3] https://github.com/zalandoresearch/fashion-mnist. The input image was resized to $32\times 32$ and the $D_{b}$ was 128. We used an Adam optimizer and the learning rate was set to 0.0002. The betas for Adam were set to (0.5, 0.999). We trained the classifier model for SFL up to 50 global training epochs and 75 local training epochs. The reconstruction model was trained up to 50 epochs under the same optimizer and settings as the classifier model. More detailed parameters are summarized at Table I.
@@
As shown in Fig. 4, increasing the depth of the cut layer (i.e., increased complexity of the client-side model) leads to higher energy consumption ($E$), causing both $E_{\text{main}}$ and $E_{\text{fed}}$ to rise due to the increased computing and networking energy usage. Concurrently, RS consistently decreases with deeper cut layers. This phenomenon implies that the depth of the cut layer can be directly correlated with the complexity of the client-side model, potentially making attackers to reconstruct the image from the smashed data more difficult. Nevertheless, due to the lack of analytic understanding of RS, a data-driven approach can be applied to model RS through offline training using empirical measurements. In this context, we adopt regression-based modeling, which is one of the most widely used approaches, including mobile CPU property models (e.g., CPU power and temperature variation modeling). Consequently, we demonstrated that a convex model can effectively approximate the behavior of RS, yielding a Root Mean Square Error (RMSE) of only $0.0028$. As a result, RS can be estimated as $\operatorname{RS}(\alpha)$ = $0.76750.3597\alpha^{2}-0.7004\alpha+$ 0.7675. It should be noted that the estimated $RS(\alpha)$ may differ across applications; therefore, it should be pre-calculated in an offline manner prior to making any adjustments. Consequently, we can optimize the selection of the cut layer, denoted by $\alpha$, to minimize privacy leakage from reconstruction attacks, as measured by $RS(\alpha)$, while ensuring that energy consumption ($E$) remains within an acceptable energy budget ($E_{\text{req}}$). Thus, the problem can be formulated as follows:
@@
[…]
@@
Figure 5: Privacy gain of proposed scheme: It represents that optimal cut layer selection ($\alpha^{*}$=0.4201) achieves minimum privacy leakage (SSIM) while satisfying the required energy budget.
III-B Other Types of Control Variables